OpenStudy (anonymous):

Water must be mixed with a 15% hydrochloric acid solution to dilute the concentration to 5%. How many liters of water should be added if we need 19 liters of the 5% solution? Round to the nearest tenth of a liter. I need an equation explained to me step by step, all the way to the end, that get's 12.7 liters as the answer. 12.7 liters is the answer given to me by my online math program.

OpenStudy (amistre64):

id try it again, but i dont really know how to explain it simpler

OpenStudy (anonymous):

Where did you get the 2.85 in your equation, in the second part...19(.05)2.85-.15w?

OpenStudy (anonymous):

you can find the liters of HCl solution: .15H=19*.05 H=6.3 then 19-6.3=12.7 12.7 liters of water.

OpenStudy (amistre64):

(19-w)(.15) .... this is just multiplication: 19 - w .15 --------- 2.85 -.15w

OpenStudy (anonymous):

I can see where you got it now...but now it's confusing...I guess I was thinking that (19-w)(.15)+w(0) was the whole equation, instead of just one part of it, amestre64. :P

OpenStudy (anonymous):

rsvitale, I kept getting 6.3 from various equations too, and would get the answer if I subtracted it from the final total of 19 liters..

OpenStudy (anonymous):

but I have no explanation or equation for why I always tended to get this answer, nor why I had to continue and subtract it from 19 to get the answer...any ideas?

OpenStudy (radar):

I would review the previous post, focus and meditate and it should become clear. amistre64 did provide a step by step (ignore the false start with w(1)) and it should become obvious even to the most casual observer. Good luck with it.

OpenStudy (anonymous):

my equation works indiana. you know how much pure HCl is in the final solution. It is 5%, so 19*.05 liters of pure HCl. you find how many liters of 15% HCl solution it takes to make 19*.05 liters of pure HCl. We use H to denote liters HCl solution: .15*H=19*.05 H=6.3 We know the rest of the total 19 litersis pure water: 19-6.3=12.7

OpenStudy (radar):

I've looked at that, but is a mixture of 6.3 Liters of Hydroclohric Acid and 12.7 liters of water a 5% solution? Looks more like a 50 % solution